the distance between two stations is 340 km.two trains start simultaneously from these stations on parallel tracks to cross each other. the speed of one of them is greater than that of the other by 5 km/hr. if the distance between the two trains after 2 hours of their start is 30 km. find the speed of each train.
Answer by ptaylor(2198)

distance(d)=rate(r) times time(t) or d=rt;t=d/r and r=d/t
Let r=rate (speed) of the first train
Then r+5=rate of the other
Distance first train travels in 2 hours =r*2 or 2r
Distance other train travels in 2 hours=(r+5)*2 or 2(r+5)
Now we know that the distance the first train travels (2r) plus the distance the other train travels (2(r+5)) equals 340-30 or 310 mi. So our equation to solve is:
2r+2(r+5)=310 mi get rid of parens
2r+2r+10=310 subtract 10 from both sides
2r+2r+10-10=310-10 collect like terms
4r=300 divide both sides by 4
r=75 mph---------------------------speed of first train
r+5=75+5= 80 mph---------------------speed of other train
CK
75*2+80*2=310
150+160=310
310=310

Hope this helps---ptaylor